# SORT3D: Spatial Object-centric Reasoning Toolbox for Zero-Shot 3D Grounding Using Large Language Models

### v1

@@ Abstract
Interpreting object-referential language and grounding objects in 3D with spatial relations and attributes is essential for robots operating alongside humans. However, this task is often challenging due to the diversity of scenes, large number of fine-grained objects, and complex free-form nature of language references. Furthermore, in the 3D domain, obtaining large amounts of natural language training data is difficult. Thus, it is important for methods to learn from little data and zero-shot generalize to new environments. To address these challenges, we propose SORT3D, an approach that utilizes rich object attributes from 2D data and merges a heuristics-based spatial reasoning toolbox with the ability of large language models (LLMs) to perform sequential reasoning. Importantly, our method does not require text-to-3D data for training and can be applied zero-shot to unseen environments. We show that SORT3D achieves state-of-the-art performance on complex view-dependent grounding tasks on two benchmarks. We also implement the pipeline to run real-time on an autonomous vehicle and demonstrate that our approach can be used for object-goal navigation on previously unseen real-world environments. All source code for the system pipeline is publicly released.  [^1] https://github.com/nzantout/SORT3D
@@ I INTRODUCTION
As we progress toward generalizable robots operating in human-centered environments like homes and offices, it is crucial for these agents to interact through natural language and align visual observations with natural language references. This capability is essential for applications such as indoor navigation robots and robot caregiving assistants. Resolving natural language expressions referring to specific objects using spatial relations—the core challenge of 3D Visual Grounding (3DVG)—remains difficult despite being an easy task for humans. For example, understanding statements such as “the chair closest to the closet door”, is a task trivial for humans [1] but still challenging for robots. While humans are usually able to identify objects from referring expressions by filtering out irrelevant objects, reasoning about spatial relationships, and utilizing semantic object attributes, such tasks remain challenging for state-of-the-art (SOTA) methods due to several reasons. First, indoor environments often contain numerous objects belonging to fine-grained classes [2], and are often specific to different environments and homes. Second, grounding complex language references to a 3D scene is difficult even for SOTA large language models (LLMs), especially when the language contains view-dependent references (e.g. “facing the door”) or use unique ways of anchoring objects that require numerical reasoning (e.g. “the second chair from the window”). Third, to train learning-based methods on such grounding tasks, a large amount of annotated data aligning language references to a 3D scene is required, which the 3D domain lacks in comparison to the 2D vision-language domain [1].
@@
While a number of existing works have developed end-to-end methods to tackle this task through training multi-modal alignment with large transformer models [3, 4, 5], these methods require large-scale annotated data, often overfitting to specific syntactic structure in training datasets while struggling to generalize to more complex utterances, resulting in only mediocre performance. More recently, numerous works have leveraged the reasoning abilities of LLMs for spatial grounding [6, 7, 8, 9, 10, 11, 12]. While some achieve strong results on structured, synthetic datasets [6], complex real-world natural language expressions such as “the tall recycling bin to the left if you are facing the door” remain challenging. These methods also often rely on careful fine-tuning or struggle with poor zero-shot performance. Additionally, such methods are not typically designed for system deployment, relying on heavily processed high fidelity reconstructions of 3D scenes [13, 14] and failing to account for constraints in model size, efficiency, and noise in the real-time semantic mapping process.
@@
To this end, we propose SORT3D, a Spatial Object-centric Reasoning Toolbox for 3D Grounding Using LLMs, shown in Fig. 2, as a novel pipeline for 3D spatial reasoning tailored towards the downstream application of object-goal navigation [15]. SORT3D is modeled intuitively after human reasoning in object disambiguation, enabling data-efficient 3D grounding without requiring annotated 3D training data. Instead, we leverage SOTA 2D vision-language models (VLMs) to extract semantic object attributes important for distinguishing objects. To efficiently filter relevant objects in large scenes with hundreds of objects, we use an open-vocabulary object filtering module, similar to how humans focus their attention to relevant objects. For complex spatial references, we utilize large language models (LLMs) with chain-of-thought [16], prompting them to reason sequentially. However, recognizing LLMs’ limitations in mathematical reasoning and spatial understanding, we introduce a modular spatial reasoning toolbox with predefined elementary spatial relations. This delegates complex spatial reasoning to structured rule-based logic, ensuring greater accuracy and interpretability. As a result, our method only requires a single in-context example of the toolbox usage and no other training data.
@@
We evaluate our method on standard 3D object referential grounding benchmarks, ReferIt3D [1] and VLA-3D [17], and show that it is competitive with SOTA performance on complex view-dependent statements while requiring less data to zero-shot generalize. We also deploy our full pipeline on a robot ground vehicle for real-time indoor navigation, demonstrating our methods ability to generalize to previously unseen, dynamic environments and run online alongside real-time perception and autonomy stacks.
@@ II RELATED WORK
Object Referential Datasets
Using referring expressions to identify a target object, also defined as the 3DVG task, has been explored in a number of datasets such as ReferIt3D [1], ScanRefer [18], SceneVerse [19], and VLA-3D [17] in the 3D domain. Of these datasets, only the Nr3D subset of ReferIt3D contains human-generated natural language utterances, while the remaining datasets generate synthetic statements using template-based generation or LLMs to rephrase. While VLA-3D generates statements synthetically, the dataset also contains mentions of both static object attributes such as color, and relative size attributes. Compared to the template-generated datasets, the human-generated statements in Nr3D contain far more complex spatial grounding language, such as “Facing the three boxes, it is the box on the right that is on top of the larger blue box.”, which requires view-point grounding, and “The lamp to the left of the desk. NOT the lamp between the beds”, which contains negation, and “the chair closest to the metal appliance” which uses coarse object references. However this is the only dataset with such reasoning challenges, indicating a lack of data for complex 3DVG and motivating the need for 3DVG methods to be data efficient while learning complex spatial reasoning.
3D Referential Object Grounding
Grounding object references to 3D scenes has been explored more extensively by various methods since the introduction of benchmarks specific to the task, including BUTD-DETR [3], MVT [4], ViL3DRel [18], 3D-VisTA [5], NS3D [7], GPS trained on SceneVerse [19], and Transcrib3D [6]. While the majority of methods utilize fusion of multi-modal information in a large transformer model [3, 4, 18, 5, 19], more recent methods have leveraged neuro-symbolic frameworks [7] and LLMs [6] to effectively reason about spatial relations. Utilizing the reasoning capabilities of LLMs in the text domain, Transcrib3D achieves overall accuracies of 70.2% on Nr3D and 98.4% on Sr3D, outperforming all previous methods after skipping statements that do not contain the target object or were answered incorrectly by human annotators. To achieve this, Transcrib3D relies on iterative code generation and fine-tuning on incorrect answers with self-reasoned corrections. While significant progress in grounding accuracy has been made, the complex spatial reasoning in the Nr3D dataset, especially statements that involve egocentric view-point, continue to pose a challenge, especially for zero-shot methods [11, 20, 9]. We aim to address this gap with our method by combining structured heuristics with the sequential reasoning capabilities of LLMs.
@@
Recent efforts have also aimed to develop 3D foundation models capable of handling general 3D tasks. These models build upon pretrained LLMs, fine-tuning them on tokenized 3D data. By leveraging the strong reasoning capabilities of LLMs, these approaches bridge the gap between natural language understanding and 3D perception. One of the pioneering works in this space, 3D-LLM [21], adopts a 2D VLM as its backbone. To extract 3D features, it projects a given 3D input into multiple 2D images, obtains dense 2D features using ConceptFusion [22], and reconstructs 3D representations from these 2D embeddings. [8] improves upon the work to handle grounding to different viewpoints and generalizes to various 3D tasks with large-scale language-scene pretraining. Another significant contribution in this domain is LEO [23], which aims to create a unified 3D foundation model. LEO utilizes a pretrained LLM and applies LoRA [24] fine-tuning to adapt it for 3D-centric tasks. A key enabling factor for these methods is their reliance on pretrained LLMs and dense 2D data, demonstrating that grounding 3D understanding in well-established 2D and language-based representations is a powerful approach for advancing 3D reasoning. This factor subsequently inspires the design of our pipeline with the incorporation of 2D captions.
Vision-Language-Navigation with LLMs
A number of recent works have also leveraged the sequential reasoning capabilities of LLMs for navigation tasks. NavGPT [25] and NavCoT [26] use VLMs to generate text descriptions of viewpoints in the scene in 2D, then task the LLM with selecting the next action in an instruction-following task. Such methods demonstrate the effectiveness of leveraging 2D visual information to guide grounding and action selection in the 3D space. While our problem setting uses object-goal navigation as a downstream task, object-centric grounding is crucial for interpreting general navigation instructions as these instructions frequently use objects to anchor viewpoints or actions.
@@ III METHODOLOGY
We formulate the 3DVG task as a precursor to downstream tasks such as object-goal navigation, instruction-following, or scene visual question answering (VQA). The input to the grounding pipeline consists of perception information from the scene and a free-form referring expression in natural language. The output is the ID of the target object referenced. Fig. 2 shows our proposed framework, which can be broken down into four components: perception of the 3D scene including a pipeline to incorporate 2D semantic information for each object, filtering for relevant objects based on the input utterance, LLM-based spatial reasoning to resolve the target object, and code generation for executing a downstream action. Each of these components is described in further detail below.
III-A Enhancing Object Perception with 2D Captions
@@
Accurately understanding the attributes and affordances of 3D objects is an essential first step for referential grounding. Existing works [6], [4] use high-level information such as object bounding boxes and labels for this task, often acquired through 3D segmentation. While 3D segmentations provide useful object-centric information, they often fail to capture fine-grained attributes like color and shape, which is a key bottle-neck of 3D object grounding models[5]. However, large-scale training for accurate 3D perception is still a challenging problem due to the lack of data and task complexity. On the other hand, VLMs trained on vast amounts of 2D data have strong priors for object understanding in 2D. Notably, VQA models excel at captioning objects in a scene. Therefore, we leverage 2D VQA models to generate descriptions for 3D objects in the scene, providing richer visual details for grounding that are otherwise missed with pure 3D perception. This approach also mirrors how humans perceive and identify objects: narrowing down candidate objects through attributes and relations to resolve ambiguities. In our approach, this finer-grained inspection is achieved by captioning cropped object images, providing a more intuitive and precise form of object grounding.
@@
A key decision to be made for captioning is what image to give the VQA model when multiple views of an object are present. We make this choice by using the viewpoint that has the highest CLIP similarity with the target label. We use the Qwen2-7B model as our VLM as we found it to work best to generate accurate and concise descriptions following our template. We query the VLM with the following prompt format: ”You are an AI model that describes the characteristics of a query object in an image. Describe the $<$object$>$ in this image, using properties like color, material, shape, affordances, and other meaningful attributes. Provide the response in this format: “The $<$object name$>$ is $<$color$>$, $<$material$>$, $<$shape$>$”. We release all object crops and captions as a supplement to the ScanNet [13] dataset along with our code. A sample of object crops and their corresponding captions are shown in Fig. 3.
@@
III-B Filtering for Relevant Objects
@@
Indoor environments such as homes can consist of hundreds of objects, few of which are relevant to a given language query or task. Thus, inspired by the ability for humans to filter out irrelevant objects and the success of past works [6, 7], we implement a filtering module that consists of: extracting referred objects along with their mentioned attributes from the input query, selecting a subset of relevant objects from the scene based on the mentioned properties, and only providing this subset of objects and properties to the downstream reasoning agent. We use the Mistral Large model [27] for these steps, filtering objects based on their text descriptions, to best leverage the ability of LLMs to process textual information.
III-C Spatial Reasoning Toolbox

### v2

@@ Abstract
Interpreting object-referential language and grounding objects in 3D with spatial relations and attributes is essential for robots operating alongside humans. However, this task is often challenging due to the diversity of scenes, large number of fine-grained objects, and complex free-form nature of language references. Furthermore, in the 3D domain, obtaining large amounts of natural language training data is difficult. Thus, it is important for methods to learn from little data and zero-shot generalize to new environments. To address these challenges, we propose SORT3D, an approach that utilizes rich object attributes from 2D data and merges a heuristics-based spatial reasoning toolbox with the ability of large language models (LLMs) to perform sequential reasoning. Importantly, our method does not require text-to-3D data for training and can be applied zero-shot to unseen environments. We show that SORT3D achieves state-of-the-art zero-shot performance on complex view-dependent grounding tasks on two benchmarks. We also implement the pipeline to run real-time on two autonomous vehicles and demonstrate that our approach can be used for object-goal navigation on previously unseen real-world environments. All source code for the system pipeline is publicly released.  [^1] https://github.com/nzantout/SORT3D
@@ I Introduction
As we progress toward generalizable robots operating in human-centered environments like homes and offices, it is crucial for these agents to interact through natural language and align visual observations with natural language references. This capability is essential for applications such as robot caregivers and indoor assistants. Resolving natural language expressions referring to specific objects using semantic object attributes and inter-object spatial relations—the core challenge of 3D referential grounding—remains difficult despite being an intuitive task for humans. For example, understanding statements such as “the chair closest to the closet door”, is a task trivial for humans [1] but still challenging for robots. While humans are usually able to identify objects from referring expressions by filtering out irrelevant objects, reasoning about spatial relationships, and utilizing semantic object attributes, such tasks remain challenging for state-of-the-art (SOTA) methods due to several reasons. First, indoor environments often contain numerous objects belonging to fine-grained classes [2], with distributions that vary widely across different homes and environments. Second, training end-to-end learning-based methods on 3D referential grounding requires a large amount of annotated data aligning language references to a 3D scene, which the 3D domain lacks in comparison to the 2D vision-language domain [1].
@@
While a number of existing works have developed end-to-end methods to tackle this task through training multi-modal alignment with large transformer models [3, 4, 5], these methods require large-scale annotated data, often overfitting to specific syntactic structure in training datasets while struggling to generalize to more complex utterances, resulting in mediocre performance. More recently, numerous works have leveraged the reasoning capabilities and rich language semantics of large language models (LLMs) for 3D referential grounding [6, 7, 8, 9, 10, 11, 12]. While some achieve strong results on benchmark datasets [6], complex real-world natural language expressions that employ both semantic attributes and spatial relations such as “the tall recycling bin to the left if you are facing the door” remain challenging. Many LLM-based approaches either struggle with poor zero-shot performance or rely on careful fine-tuning or heavy prompt engineering tailored towards benchmark datasets. Consequently, such methods are not typically designed for system deployment, relying on high fidelity reconstructed meshes of 3D scenes [13, 14] and failing to account for constraints in model size, efficiency, and noise in the real-time semantic mapping process.
@@
To this end, we propose SORT3D, a Spatial Object-centric Reasoning Toolbox for 3D Grounding Using LLMs, shown in Figure 2, as a novel pipeline for 3D spatial reasoning tailored towards the downstream application of object-goal navigation [15]. SORT3D is modeled intuitively after human reasoning in object disambiguation, enabling data-efficient 3D referential grounding without requiring annotated 3D training data. To achieve this, we decompose the task into a three-stage approach. After obtaining object names and bounding boxes using an instance-level semantic mapping module, we leverage SOTA 2D vision-language models (VLMs) to extract semantic object attributes important for distinguishing objects. We then use an open-vocabulary object filtering module to efficiently filter relevant objects in large scenes with hundreds of objects, similar to how humans focus their attention to relevant objects. We finally leverage the strong semantic language and sequential reasoning priors of LLMs using chain-of-thought prompting [16] to parse a complex referential statement into a series of function calls to our spatial reasoning toolbox, containing search functions that return the IDs of objects that satisfy heuristically defined elementary spatial relations. This delegates complex relative spatial reasoning to structured rule-based logic, ensuring greater accuracy and interpretability. As a result, our method only requires a single in-context example of the toolbox usage and no other training data.
@@
We evaluate our method on standard 3D object referential grounding benchmarks, ReferIt3D [1] and IRef-VLA [17], and demonstrate performance competitive with SOTA on complex view-dependent statements while requiring less data to zero-shot generalize. We also deploy our full pipeline on two robotic ground vehicles for real-time indoor navigation, demonstrating our method’s ability to further generalize to previously unseen environments and run online alongside real-time perception and autonomy stacks.
@@ II Related Work
Referential Object Grounding Datasets
Using referring expressions to identify a target object in a 3D scene, defined as the 3D referential grounding task, has been explored in a number of datasets such as ReferIt3D [1], ScanRefer [18], SceneVerse [19], and IRef-VLA [17]. Of these datasets, only the Nr3D subset of ReferIt3D and ScanRefer contain human-generated natural language utterances, while the remaining datasets generate synthetic statements using template-based generation and/or LLMs to rephrase. Synthetic datasets only employ basic spatial relations like ”lamp by the desk” in referential statements, although IRef-VLA adds the usage basic semantic object attributes—color and size—to create unambiguous references. Compared to the template-generated datasets, the human-generated statements in Nr3D contain far more complex spatial grounding language, such as “Facing the three boxes, it is the box on the right that is on top of the larger blue box.”, which requires view-point grounding, “The lamp to the left of the desk. NOT the lamp between the beds”, which contains negation, and “the chair closest to the metal appliance” which uses coarse object references and semantic attribute. This complexity, coupled with the small scale of human-generated 3D referential language data, motivates the need for 3D referential grounding methods to leverage strong semantic language priors from other sources (like LLMs), and use a more structured approach for grounding.
3D Referential Object Grounding Baselines
Grounding object references to 3D scenes has been explored more extensively by various methods since the introduction of benchmarks specific to the task. Approaches to this task include end-to-end models like BUTD-DETR [3], MVT [4], ViL3DRel [20], 3D-VisTA [5], and GPS trained on SceneVerse [19], which fuse multi-modal information in large transformer models and are trained and fine-tuned directly on referential grounding benchmarks. More recent methods decompose the task, leveraging neuro-symbolic frameworks like NS3D [7], and LLMs and VLMs in a zero-shot manner to effectively reason about spatial relations like ZSVG3D [21], VLM-Grounder [11], CSVG [9], and Transcrib3D [6]. Utilizing the reasoning capabilities of LLMs in the text domain, Transcrib3D achieves overall accuracies of 70.2% and 98.4% respectively on subsets of Nr3D and Sr3D, outperforming all previous methods. To achieve this, Transcrib3D represents a scene as a list of object names, colors, and positions, and relies on iterative code generation and benchmark-specific prompt engineering giving the LLM guiding principles for grounding. Transcrib3D additionally fine-tunes smaller LLMs on incorrect answers with corrections self-reasoned by larger LLMs. While significant progress in grounding accuracy has been made, the complex spatial reasoning required for the Nr3D dataset, especially with statements that involve egocentric viewpoints and utilize object semantics, continues to pose a challenge, especially for zero-shot methods [11, 21, 9, 6]. We aim to address this gap with our method by combining structured heuristics and vision foundation models with the sequential reasoning capabilities of LLMs.
@@
Recent efforts have also aimed to develop 3D foundation models capable of handling general 3D tasks. These models leverage the strong reasoning capabilities of pretrained LLMs, achieving strong performance on 3D scene understanding tasks by fine-tuning the LLMs on tokenized 3D data. One of the pioneering works in this space, 3D-LLM [22], distills rich 2D foundation model features and a 2D VLM backbone to improve performance, mitigating the lack of 3D data. [8] improves upon the work to handle grounding to different viewpoints and generalizes to various 3D tasks with large-scale language-scene pretraining. The efficacy of 2D foundation model features in these approaches demonstrates that grounding 3D understanding in well-established 2D and language-based representations is a powerful approach for advancing 3D reasoning. This factor subsequently inspires the design of our pipeline with the incorporation of 2D captions.
Vision-and-Language Navigation with LLMs
A number of recent works have also leveraged the sequential reasoning capabilities of LLMs for navigation tasks. NavGPT [23] and NavCoT [24] use VLMs to generate text descriptions of viewpoints in the scene in 2D, then task the LLM with selecting the next action in an instruction-following task. Such methods demonstrate the effectiveness of leveraging 2D visual information to guide grounding and action selection in the 3D space. However, within a collaborative setting between a human and a robotic agent, the human would more commonly use single referential statements like ”fetch the tv remote on the cabinet” and assume the robot has full knowledge the scene rather than describing a full trajectory towards an object. For building a 3D referential grounding-based VLN system with object-goal navigation as the basic downstream task, text descriptions of landmarks must also be combined with a structured map representing the scene.
@@ III Methodology
In a human-agent collaborative setting, humans commonly refer to objects in the scene by relating them to other objects using commands like ”grab the red mug in the top left cupboard over the sink”. Finding the correct object being referenced from the utterance is the task of 3D referential grounding, which is fundamental for the deployment of a practical VLN system. 3D referential grounding additionally acts as a precursor to downstream tasks such as object-goal navigation, multi-action instruction-following, and scene visual question answering (VQA). We therefore present SORT3D, a zero-shot pipeline for 3D referential grounding, which decomposes the task into multiple subtasks, leverages foundation models to obtain robust zero-shot performance, and targets downstream mobile robot navigation in a real-world environment.
@@
The input to the grounding pipeline consists of perception information from the scene and a free-form referring expression in natural language. The output is the ID of the target object referenced. Figure 2 shows our proposed framework, which can be broken down into four components: (i) an instance-level semantic mapping system to obtain 3D bounding boxes for real-world deployment, (ii) a captioning pipeline to incorporate rich 2D semantic information for each object, (iii) filtering for relevant objects based on the input utterance, and (iv) LLM-based reasoning augmented with a spatial reasoning toolbox to resolve the target object, followed by code generation for executing a downstream action. Each of these components is described in further detail below.
III-A Instance-level Semantic Mapping
For our real-world experiments, we use an object instance-level semantic mapping module running in real-time to obtain the 3D bounding boxes to be input into the LLM and the spatial reasoning toolbox. This component is the only pipeine change required for our method to be deployed in the real-world. Our mobile robot perception setup for real-world experiments consists of a 360 camera and a 3D LiDAR (section V-C contains further hardware details). We initially perform object detection in 2D using open-vocabulary object detection [25] and instance segmentation [26] models. We then project the registered LiDAR point clouds onto the semantic images and associate each point with its corresponding pixel semantic ID. As the robot moves and produces new observations, we associate per-frame object instance pointclouds using a 2D tracking module and 3D proximity priors, followed by filtering steps to obtain 3D instance pointclouds [^2] For further details on the semantic mapping module, see https://github.com/gfchen01/semantic_mapping_with_360_camera_and_3d_lidar. The usage of open-vocabulary 2D foundation models allows our semantic mapping module to generalize to new environments as we show in our real-world experiments (section V-C). For our results on the ReferIt3D benchmarks, we simply use ground truth bounding boxes and instance segmentations.
III-B Enhancing Object Perception with 2D Captions
@@
Accurately understanding the attributes and affordances of 3D objects is an essential first step for referential grounding. Existing works [6], [4] use high-level information such as object bounding boxes and labels for this task, often acquired through 3D segmentation. While 3D segmentation provides useful object-centric information, it often fails to capture fine-grained attributes like color and shape, which is a key bottleneck of 3D object grounding models [5]. Nonetheless, large-scale training for accurate 3D perception is still a challenging problem due to the lack of data and task complexity. On the other hand, VLMs trained on vast amounts of 2D data have strong priors for object understanding in 2D. Notably, VQA models excel at captioning objects in a scene. Therefore, we leverage 2D VQA models to generate descriptions for 3D objects in the scene, providing richer visual details for grounding that are otherwise missed with pure 3D perception. This approach also mirrors how humans perceive and identify objects: narrowing down candidate objects through attributes and relations to resolve ambiguities. In our approach, this finer-grained inspection is achieved by generating captions for each object from cropped object images, providing a more intuitive and precise form of object grounding.
@@
A key decision to be made for captioning is what image to give the VQA model when multiple views of an object are present. We make this choice by using the viewpoint that has the highest CLIP similarity with the target label, which is obtained from ground truth for the ReferIt3D benchmarks and from the semantic mapping module for the real-world experiments. We use Qwen2-VL-7B [27] as our VLM as we found it to perform best in generating accurate and concise descriptions following our template, and the quantized version of Qwen2.5-VL-Instruct-3B for system deployment due to memory constraints. We query the VLM with the following prompt format: ‘‘You are an AI model that describes the characteristics of a query object in an image. Describe the <<object>> in this image, using properties like color, material, shape, affordances, and other meaningful attributes. Provide the response in this format: ‘The <<object name>> is <<color>>, <<material>>, <<shape>>’". We release all object crops and captions as a supplement to the ScanNet [13] dataset along with our code. A sample of object crops and their corresponding captions are shown in Figure 3.
@@
III-C Filtering for Relevant Objects
@@
Indoor environments such as homes can consist of hundreds of objects, of which only a few are relevant to a given language query or task. Thus, inspired by the ability for humans to filter out irrelevant objects and the success of past works [6, 7], we implement an LLM-based filtering module consisting of two queries. Given an input command like “The nightstand to the right of the bed”, the first query extracts object nouns and modifiers (e.g. nightstand and bed), and the second query returns the relevant object instances from the list of IDs and names provided by the perception module (in this case returning all nightstands and beds in the scene). We use Mistral Large 2 [28] for these steps, filtering objects based on their text descriptions, to best leverage the ability of LLMs to process textual information.
III-D Spatial Reasoning Toolbox